What is the t statistic for the following two samples?

  • Sample 1:
    • n=4
    • SS = 1.94
    • M = 2.5
  • Sample 2
    • n=5
    • SS = 3.96
    • M = 4.0
Answer #1

Given:

· Sample 1:

o n=4

o SS = 1.94

o M = 2.5

· Sample 2

o n=5

o SS = 3.96

o M = 4.0

S1^2 = SS/(n – 1) = 1.94/(4 – 1) = 0.646667

S2^2 = SS/(n – 1) = 3.96/(5 – 1) = 0.99

S1 = sqrt(0.646667) = 0.804156

S2 = sqrt(0.99) = 0.994987

The formula for t statistic is given as below:

t = (X1bar – X2bar) / sqrt[(S12 / n1)+(S22 / n2)]

t = (2.5 – 4.0) / sqrt[(0.646667/4) + (0.99/5)]

t = -1.5/0.599722

t = -2.50116

t statistic = -2.50116